# The Fast and the Frame-Dragging: Efficient waveforms for asymmetric-mass eccentric equatorial inspirals into rapidly-spinning black holes

### v1

@@ V.3 Parameter recovery with approximate mode amplitude models
At each gridpoint, inference of the true signal with the approximate model generically introduces systematic biases in the best-fit parameter estimate that can be approximated in the high-SNR limit by the linear signal approximation given by Cutler and Vallisneri Cutler:2007mi Eq. 60). To quantify these biases, we calculate the sigma-contour level between the injected and the best-fit points, given by the Mahalanobis distance mahalanobis , notated as $D_{\rm Maha}$ and defined explicitly in Eq. 63 of Appendix D. $D_{\rm Maha}$ quantifies how many sigma-levels away the biased parameter recovery point is from the injected parameters in the full $D$-dimensional space. It scales proportionately with the induced biases on the recovered parameters, such that $D_{\rm Maha}\in[0.0,\infty)$ and the null value is recovered at the injected parameter point. See discussion below Eq. 63 for more details. In the rest of this section, we use “Mahalanobis distance” and “sigma contours” interchangeably. We present our main results in the following two subsections. The first subsection (results \@slowromancapi@) quantifies the sigma contour levels in the full-dimensional and the marginalized 1-dimensional spaces across the $(a,e_{0})$ grid for a fixed total mass binary, while the second subsection (results \@slowromancapii@) assesses the impact that different total EMRI masses ($m_{1}+m_{2}$) have on the biases across the grid. We provide qualitative reasoning for the observed trends in both subsections and draw our conclusions in the final paragraph.
V.3.1 Results \@slowromancapi@: Sigma contour levels
Our results are presented in Fig. 16 for the $(m_{1},m_{2})=(10^{6},10)M_{\odot}$ case. For the $\ell$2 case (top panel), we find that the sigma contours categorically scale with both $a$ and $e_{0}$. On the contrary, in the case of AAK (bottom panel), we observe an inverse scaling with $e_{0}$, with the typical sigma-contour levels $\sim 8-10$ times higher than $\ell$2. We now qualitatively explain these scalings.
Scaling with $a$—At larger $a$’s, the separatrix of the orbit is closer to the horizon of the MBH, which will increase the number of highly-relativistic orbits. In this region, the importance of higher-order and fully relativistic mode amplitudes increases, and consequently, parameter recovery with the $\ell$2 and AAK models is worsened.
Scaling with $e_{0}$— For the $\ell$2 model, as the eccentricity increases, higher $\ell$-mode amplitudes are no longer suppressed in general, and are more similar in magnitude to the $\ell=2$ mode (cf. Refs. Drasco:2005kz ; Hughes:2021exa ; Fujita:2009us .) This leads to worsened parameter recovery with the $\ell$2 model. On the other hand, the AAK model does not scale as strongly with eccentricities, neither is the scaling monotonic across the grid. While this trend is qualitatively consistent with the findings in Fig. 15 (see also Section V.2), our results warrant a detailed analysis in the future. We argue that at smaller initial eccentricities where the separatrix is closer to the MBH, the CO completes more strong-field orbits, and thus the AAK model incurs larger biases.
We also calculated the 1-dimensional marginalized sigma-biases ($z$-scores) on each parameter explicitly and found that the luminosity distance $d_{\mathrm{L}}$ incurs the largest biases among all parameters with a median value $\sim 0.3\sigma$ for $\ell$2 and $\sim 1.5\sigma$ for AAK when $(m_{1},m_{2})=(10^{6},10)\,M_{\odot}$. In both cases, the shift in $d_{\mathrm{L}}$ can be understood to make up for the loss of SNR compared to the true signal when it is recovered using the approximate waveforms (also see Section V.2). In the $\ell$2 model, all other parameters are typically biased to $\lesssim 0.1\,\sigma$, showing robust recovery. However, the biases in the AAK model are a factor $\sim 3-5$ times higher (but still within $1\,\sigma$), consistent with the trend in Fig. 16. Two notable exceptions are the recovered phases $(\Phi_{\phi_{0}},\Phi_{r_{0}})$, with biases a factor $\sim 10-12$ higher compared to the $\ell$2 model, hinting at severe dephasing of the AAK waveform compared to the injected signal.
V.3.2 Results \@slowromancapii@: Other mass pairs
Finally, we analysed two additional EMRI sources with (i) $(m_{1},m_{2})=(0.5\times 10^{6},5)\,M_{\odot}$ (lower-mass), and (ii) $(m_{1},m_{2})=(1.5\times 10^{6},15)\,M_{\odot}$ (higher-mass). In the $\ell$2 case, we found the qualitative scaling of sigma contours with $a$ and $e_{0}$ to be the same. However, for AAK, the scaling with $e_{0}$ significantly depended on the total mass of the system: biases incurred by the lower-mass EMRI were largely insensitive to $e_{0}$ and scaled with $a$, while the higher-mass EMRI showed stronger dependence on both $a$ and $e_{0}$. While a detailed analysis is beyond the scope of this work, these scalings can be qualitatively attributed to the relation between the number of relativistic orbits completed, the total mass of the system, and the LISA sensitivity curve, similarly to the discussion in Section V.2.
@@ V.3.2 Results \@slowromancapii@: Other mass pairs
AAK case—The typical biases incurred by the lower-mass EMRI were the largest in this case, opposite to the trend observed in $\ell$2. This may be attributed to the larger number of strong-field orbits completed by the lower-mass EMRI, leading to a larger accumulation of phase errors. The marginalized 1-dimensional sigma biases showed that, while other parameters typically incurred similar biases across all three systems, the lower-mass binary incurred $\approx 40\%$ larger biases in the initial phases $(\Phi_{\phi_{0}},\Phi_{r_{0}})$ than the higher-mass system. This is consistent with our interpretation.
Overall, we find that the biases incurred using approximate models are small across a broad grid of $a$ and $e_{0}$ values. Even though the AAK waveforms led to sigma contour level biases of factor $8-10$ larger than $\ell$2, both models were able to recover all inferred parameters with 1-dimensional biases at $\lesssim 1\,\sigma$. This is useful, e.g., for coarse-grain recovery of the parameters during the search stage of EMRI inference, for which faster and approximate waveforms may be necessary and sufficient gair2004event . We caution, however, that inference with approximate waveforms may lead to multimodalities in the likelihood surface katz2021fast ; chua2022nonlocal , potentially compromising robust recovery. The linear signal approximation framework adopted in our study does not capture such features, and we leave its detailed analysis to future work. Additionally, the absence of accurate models may significantly bias the inference of fully relativistic signals emitted from systems with high primary spins. The accuracy of parameter recovery also strongly depends on the frequency band of the signal, the orbital eccentricity, and the choice of the approximate model.
@@ V.4 Horizon redshift of asymmetric-mass binaries with LISA
LISA is expected to be sensitive to a wide range of GW signals emitted by the inspiral of both stellar-mass and intermediate-mass COs into MBHs. Depending on the mass ratios of these sources, observations of IMRIs and EMRIs will be effective probes of the origin and evolution of MBHs and their environments in mass ranges currently inaccessible via electromagnetic means imbh ; LISA:2024hlh . In this subsection, we will investigate the range of distances (and therefore how far back in cosmic history) over which LISA can detect these sources. Figure 17 shows sky-averaged horizon redshifts $\bar{z}$ for asymmetric-mass binaries of various component masses, MBH spins and initial eccentricities. Each line consists of $20$ points, with detector-frame primary masses spaced uniformly in their logarithm between $5\times 10^{4}$ and $5\times 10^{8}\,M_{\odot}$. We restrict the two panels in Fig. 17 to the region of primary source mass $m_{1}^{\mathrm{(s)}}\in\left[10^{5},10^{7}\right]\,M_{\odot}$.
@@
For each combination of intrinsic parameters, we draw $100$ sets of extrinsic parameters (initial phases and orientation / sky-position angles) according to Appendix C, and average the SNRs of the corresponding systems. We compute these SNRs according to the conventions outlined in Appendix D, including the unresolvable Galactic foreground in the noise PSD.
We truncate the frequency band of each SNR computation according to the largest harmonic frequency $\omega_{mn}^{(\mathrm{max})}$ one hour before plunge. Interpolation errors in the application of the LISA response (which enter in the construction of the projections and the time-delay interferometry (TDI) delays) manifest as a broad high-frequency spectral feature; as the contribution of the signal to the spectrum tapers strongly above $\omega_{mn}^{(\mathrm{max})}$ (accounting for leakage effects), the impact of these errors can be significantly reduced by truncating the inner product at this frequency. This effect is especially prominent in high-mass and high-mass ratio systems, and the masking procedure we implement here will significantly improve the stability of parameter estimation studies for these sources.
@@
We then compute the luminosity distance $d_{\mathrm{L}}$ at which each system should be placed to achieve an SNR detection threshold of 20 (a value commonly assumed in the literature Bonetti:2020jku ; Pozzoli:2023kxy ; chua2017augmented ) and convert it to $\bar{z}$ assuming Planck18 cosmology Planck:2018vyg ; Lahav:2024npe . Finally, to take into account the dependence on the number of generated sources during the averaging procedure and the discreteness of the mass grid used for the data generation, we fit a Gaussian Process Regressor (GPR) gpr ; scikit-learn on the computed $(m_{1}^{\mathrm{(s)}},\bar{z})$ points. This allows us to provide an estimate of the uncertainty associated with each horizon redshift curve, and to extrapolate to regions beyond the parameter space coverage of our model for two-year inspirals. This extrapolation applies for large masses, large spins, and high eccentricities (which correspond to highly eccentric inspirals at low orbital separations), and leads to the large uncertainty associated with the high-mass tail of the $a=0.999$ curve in the right-hand panel of Fig. 17.
@@
We find that the horizon redshift only mildly depends on the initial eccentricity for all values of $a$ explorable with our waveform model. The left panel of Fig. 17 shows the mass-ratio dependence of the horizon redshift for two different values of $e_{0}$, namely $0.5$ (higher eccentricity; solid lines) and $0.1$ (lower eccentricity; dashed lines). We find that the maximum horizon redshift of the lower-eccentricity curve is only slightly smaller than that of the higher-eccentricity curve, with the relative difference between the maxima of the two eccentricity curves being less than $6\%$ across all mass ratios considered. More interestingly, the two show different correlations with the primary mass; in particular, the lower-eccentricity curves do not seem to show the SNR “hump” we see in the higher-eccentricity curves. This is especially evident in the $\epsilon\gtrsim 10^{-3}$ cases, and can be interpreted as the contribution of higher-frequency waveform modes with large $n$; at larger total masses, these modes lie in the most sensitive region of the LISA sensitivity curve (shown in Fig. 1) and contribute significantly to the total SNR of the system. As the amplitudes of these modes scale strongly with eccentricity, this feature is more readily apparent for particularly eccentric systems. Additionally, the value of $m_{1}^{\mathrm{(s)}}$ at which $\bar{z}$ is maximised varies with respect to $e_{0}$. The deviation is reasonably small, with a relative difference between the two eccentricity cases (over all primary mass values) of $\sim 18\%$ at its largest. This effect depends on the mass ratio: the low-eccentricity peak is shifted toward smaller masses for the $\epsilon=10^{-6}$ systems, while it is moved toward larger values for the remaining systems considered, with larger shifts for larger mass ratios. The most significant contribution to the SNR occurs as the CO enters the strong field (towards the end of inspiral); due to our initial conditions on $p_{0}$ in this analysis being fixed by the time-to-plunge, the eccentricity at this stage of the inspiral decreases as a function of mass ratio (as the inspiral will circularise more when starting from larger orbital separations). This accentuates the impact of eccentricity on detectability as mass ratio increases, as depicted in the left panel of Fig. 17.
@@
We also find that the mass location of the detectability peak shows different trends in the two panels, moving towards lower (higher) source frame primary masses with larger mass-ratios (MBH spins). The relationship between mass ratio and SNR (left panel), for a given value of $m_{1}$, is largely driven by the proportionality of the waveform amplitudes to $\mu\sim m_{2}$ (Eq. 1). As the rate of inspiral also increases with mass ratio, our fixed time-to-plunge leads to higher values of $p_{0}$; as this decreases the frequencies of strong harmonics, this in turn slightly accentuates the detectability of lower-mass systems, which manifests as a gradual shift in the location of the maximum horizon redshift in the left panel of Fig. 17. The impact of MBH spin (right panel) is also highly pronounced, particularly at larger values of $m_{1}$. Raising MBH spin increases the number of strong-field orbital cycles completed by the CO before it plunges, where both the amplitude and frequency of the GWs emitted are largest. For lower-$m_{1}$ systems, this high-frequency radiation sits towards the upper end of LISA’s sensitivity and therefore does not contribute strongly to the SNR (and therefore the detectability) of these systems. As $m_{1}$ increases, this strong-field emission decreases towards the frequencies where LISA is most sensitive, significantly enhancing the overall SNR. This trend continues until the strong-field emission shifts below the minimum in the LISA sensitivity curve ($\sim 3\,\mathrm{mHz}$; see Fig. 1) at which point the SNR begins to decrease.
@@ V.5 Distinguishing quasi-circular and mildly eccentric systems
Figure 18: Mismatches (blue dots) between quasi-circular ($h^{\mathrm{circ}}$) and eccentric $h^{\mathrm{ecc}}$ waveforms as $e_{0}\rightarrow 0$. Other source parameters $\{m_{1},m_{2},a\}=\{10^{6}\,M_{\odot},25\,M_{\odot},0.998\}$, with SNR $\rho=100$. The black dashed line indicates the Lindblom distinguishability criterion (Eq. 15) for this SNR, with $D=12$ (12 sampled parameters). For $e_{0}\gtrsim 10^{-5}$, mismatches grow as $(e_{0})^{4}$ (red dashed line) in line with PN scaling arguments. At lower initial eccentricities, other systematics in waveform generation obfuscate any physical relationship between mismatch and initial eccentricity.
@@
We compute mismatches $\mathcal{M}(h^{\rm circ},h^{\rm ecc})$ for $e_{0}\in[10^{-8},10^{-3}]$. Our results are given in Fig. 18. For very low eccentricities $e_{0}\in[10^{-8},10^{-5})$, the mismatch between circular and eccentric waveforms is dominated by ODE numerical errors. However, for $e_{0}\gtrsim 10^{-5}$, the mismatch grows by an amount proportional to $(e_{0}^{2})^{2}$ as given by the red dashed line. The black dashed line indicates the Lindblom distinguishability criterion (Eq. 15), where mismatch values $(<)>D/2\rho^{2}$ indicate that the true and approximate waveforms are (in)distinguishable. As in Section IV.3, we set the number of sampled model parameters $D=12$. From Fig. 18, we see that the two models are distinguishable for $e_{0}\gtrsim 2\cdot 10^{-4}$, highlighting that for $\rho<100$ we cannot resolve eccentricities smaller than $e_{0}<2\cdot 10^{-4}$ for this set of source parameters. We have checked that our conclusions remain unchanged when decreasing the ODE integrator error $\sigma_{\mathrm{tol}}$.
@@ V.6 Inference prospects for asymmetric-mass binaries with LISA
$1$ EMRI (Prograde) $10^{6}$ $10^{1}$ $0.998$ $7.728$ $0.730\,(0.045)$ $2.204$ $0.394$ $50$
$2$ Strong-field EMRI $10^{7}$ $10^{1}$ $0.998$ $2.120$ $0.425\,(0.261)$ $3.590$ $0.593$ $30$
$3$ Heavy IMRI $10^{7}$ $10^{5}$ $0.950$ $23.425$ $0.850\,(0.023)$ $7.250$ $1.058$ $500$
$4$ Light IMRI $10^{5}$ $10^{3}$ $0.950$ $74.383$ $0.850\,(0.004)$ $3.500$ $0.581$ $200$
$5$ EMRI (Retrograde) $10^{5}$ $10^{1}$ $-0.500$ $26.190$ $0.800\,(0.195)$ $1.567$ $0.294$ $30$
@@
Table 1: Parameters for investigated EMRI and IMRI sources. Waveforms for these systems are shown in Fig. 19. (From left to right columns): Detector-frame masses $m_{1}$ and $m_{2}$, dimensionless MBH spin $a$, initial semi-latus rectum $p_{0}$, initial (final) eccentricity $e_{0}$ ($e_{\mathrm{f}}$), luminosity distance $d_{\mathrm{L}}$, redshift $z_{\mathrm{r}}$ and the optimal SNR $\rho$. Each source is observed for two years and plunges just after the end of the observation window. For sources with $m_{1}=10^{5}M_{\odot}$, we use a finer sample cadence $\mathrm{d}t=2\,\mathrm{s}$ to resolve high-frequency GWs near plunge; all other sources use our fiducial sampling interval of $\mathrm{d}t=5\,\mathrm{s}$. In all cases, we fix the angular parameters $(\theta_{S},\phi_{S},\theta_{K},\phi_{K})=(0.8,2.2,1.6,1.2)$ and initial orbital phases $(\Phi_{\phi_{0}},\Phi_{r_{0}})=(2.0,3.0)$.
@@
In this subsection, we investigate parameter estimation prospects for a range of IMRI and EMRI sources. As few is capable of generating waveforms in $\sim 100\,\mathrm{ms}$ (Section V.1), we are able to perform Bayesian inference studies on a timescale of hours, directly exploring the structure and scale of the posterior distributions of these sources. The motivation for performing these analyses is twofold. First, we seek to verify that our model reproduces features expected of posterior distributions for these systems: in particular, that they are approximately Gaussian, and that parameters intrinsic to the inspiral dynamics (such as spin and eccentricity) are recovered with high precision. Second, it was identified in Ref. Katz:2021yft that the mode content of EMRI waveform models can be reduced significantly without significantly biasing inference results; our analyses will re-examine this conclusion with the inclusion of MBH spin and for IMRI sources. To our knowledge, this is the first full Bayesian investigation of rapidly-spinning and highly eccentric IMRI and EMRI systems to appear in the literature to date (however, see Ref. Huez:2025npe for a parameter estimation study of mildly eccentric binaries with $q\leq 10$).
@@
Figure 19: Time-domain snapshots of the science-case waveforms described in Table 1 during the early (left panel) and late (right panel) periods of a two-year inspiral. This set of representative sources demonstrates the wide range of waveform morphologies captured by our model. In all the panels, only the plus polarization $h_{+}$ is shown; for the two IMRI systems, initial orbital phases have also been adjusted for visual clarity.
@@
Following the data analysis conventions and methods outlined in Appendix D, we perform parameter estimation on simulations of each of the sources listed in Table 1. The parameters of these sources were chosen to represent the wide range of waveform morphologies accessible with our waveform model Fig. 19, with luminosity distances set to astrophysically-motivated values. A notable exception to our analysis conventions is that we set a finer sampling cadence of $\mathrm{d}t=2\,\mathrm{s}$ for Sources 4 and 5 in order to better resolve their high-frequency mode content; for other sources, we use our fiducial sampling cadence of $\mathrm{d}t=5\,\mathrm{s}$. For each simulation, we generate a waveform with mode-selection parameter $\kappa=0$ (i.e., all modes in our amplitude model are included). During inference, we consider $\kappa=10^{-5}$ and $\kappa=\{10^{-3},10^{-2}\}$ (the latter value depending on the SNR of the injection), demonstrating how parameter inference results are impacted when weaker harmonics (which are more important for high eccentricities and spins; see Fig. 4) are neglected. We do not show the (marginal) posteriors obtained from our inference runs in the main text for brevity; they can be found in Appendix F, and are referenced in the right-most column of Tab. 2, where we summarize the relative measurement precisions for parameters of astrophysical interest. These include the source frame masses $m_{i}^{(\rm s)}$, spin $a$, final eccentricity $e_{f}$ and luminosity distance $d_{\mathrm{L}}$. We also report the sky-localization area of each source: given the covariance matrix of the sky-angle posterior samples $\Sigma=\rm cov(\theta_{S},\,\phi_{S})$, and $\bar{\theta}_{S}=\rm{median}(\theta_{S})$, we compute the $99\%$ quantile of the sky-localization area in square degrees, given by

### v2

@@ V.3 Parameter recovery with approximate mode amplitude models
At each gridpoint, inference of the true signal with the approximate model generically introduces systematic biases in the best-fit parameter estimate that can be approximated in the high-SNR limit by the linear signal approximation given by Cutler and Vallisneri Cutler:2007mi ( Eq.˜61). To quantify these biases, we calculate the sigma-contour level between the injected and the best-fit points, given by the Mahalanobis distance mahalanobis , notated as $D_{\rm Maha}$ and defined explicitly in Eq.˜64. $D_{\rm Maha}$ quantifies how many sigma-levels away the biased parameter recovery point is from the injected parameters in the full $D$-dimensional space. It scales proportionately with the induced biases on the recovered parameters, such that $D_{\rm Maha}\in[0.0,\infty)$ and the null value is recovered at the injected parameter point. See discussion below Eq.˜64 for more details. In the rest of this section, we use “Mahalanobis distance” and “sigma contours” interchangeably. We present our main results in the following two subsections. The first subsection (results I) quantifies the sigma contour levels in the full-dimensional and the marginalized 1-dimensional spaces across the $(a,e_{0})$ grid for a fixed total mass binary, while the second subsection (results II) assesses the impact that different total EMRI masses ($m_{1}+m_{2}$) have on the biases across the grid. We provide qualitative reasoning for the observed trends in both subsections and draw our conclusions in the final paragraph.
V.3.1 Results I: Sigma contour levels
[…]
V.3.2 Results II: Other mass pairs
[…]
@@ V.3.2 Results II: Other mass pairs
[…]
@@ V.4 Horizon redshift of asymmetric-mass binaries with LISA
LISA is expected to be sensitive to a wide range of GW signals emitted by the inspiral of both stellar-mass and intermediate-mass COs into MBHs. Depending on the mass ratios of these sources, observations of IMRIs and EMRIs will be effective probes of the origin and evolution of MBHs and their environments in mass ranges currently inaccessible via electromagnetic means imbh ; LISA:2024hlh . In this subsection, we will investigate the range of distances (and therefore how far back in cosmic history) over which LISA can detect these sources. Figure˜17 shows sky-averaged horizon redshifts $\bar{z}$ for asymmetric-mass binaries of various component masses, MBH spins and initial eccentricities. Each line consists of $20$ points, with detector-frame primary masses spaced uniformly in their logarithm between $5\times 10^{4}$ and $5\times 10^{8}\,M_{\odot}$. We restrict the two panels in Fig.˜17 to the region of primary source mass $m_{1}^{\mathrm{(s)}}\in\left[10^{5},10^{7}\right]\,M_{\odot}$.
@@
For each combination of intrinsic parameters, we draw $100$ sets of extrinsic parameters (initial phases and orientation / sky-position angles) according to Appendix˜C, and average the SNRs of the corresponding systems.
@@
We then compute the luminosity distance $d_{\mathrm{L}}$ at which each system should be placed to achieve an SNR detection threshold of 20 (a value commonly assumed in the literature Bonetti:2020jku ; Pozzoli:2023kxy ; chua2017augmented ) and convert it to $\bar{z}$ assuming Planck18 cosmology Planck:2018vyg ; Lahav:2024npe . Finally, to take into account the dependence on the number of generated sources during the averaging procedure and the discreteness of the mass grid used for the data generation, we fit a Gaussian Process Regressor (GPR) gpr ; scikit-learn on the computed $(m_{1}^{\mathrm{(s)}},\bar{z})$ points. This allows us to provide an estimate of the uncertainty associated with each horizon redshift curve, and to extrapolate to regions beyond the parameter space coverage of our model for two-year inspirals. This extrapolation applies for large masses, large spins, and high eccentricities (which correspond to highly eccentric inspirals at low orbital separations, which are outside of the domain of validity of our model), and leads to the large uncertainty associated with the high-mass tail of the $a=0.999$ curve in the right-hand panel of Fig.˜17.
@@
We find that the horizon redshift only mildly depends on the initial eccentricity for all values of $a$ explorable with our waveform model. The left panel of Fig.˜17 shows the mass-ratio dependence of the horizon redshift for two different values of $e_{0}$, namely $0.5$ (higher eccentricity; solid lines) and $0.1$ (lower eccentricity; dashed lines). We find that the maximum horizon redshift of the lower-eccentricity curve is only slightly smaller than that of the higher-eccentricity curve, with the relative difference between the maxima of the two eccentricity curves being less than $6\%$ across all mass ratios considered. More interestingly, the two show different correlations with the primary mass; in particular, the lower-eccentricity curves do not seem to show the SNR “hump” we see in the higher-eccentricity curves. This is especially evident in the $\epsilon\gtrsim 10^{-3}$ cases, and can be interpreted as the contribution of higher-frequency waveform modes with large $n$; at larger total masses, these modes lie in the most sensitive region of the LISA sensitivity curve (shown in Fig.˜1) and contribute significantly to the total SNR of the system. As the amplitudes of these modes scale strongly with eccentricity, this feature is more readily apparent for particularly eccentric systems. Additionally, the value of $m_{1}^{\mathrm{(s)}}$ at which $\bar{z}$ is maximised varies with respect to $e_{0}$. The deviation is reasonably small, with a relative difference between the two eccentricity cases (over all primary mass values) of $\sim 18\%$ at its largest. This effect depends on the mass ratio: the low-eccentricity peak is shifted toward smaller masses for the $\epsilon=10^{-6}$ systems, while it is moved toward larger values for the remaining systems considered, with larger shifts for larger mass ratios. The most significant contribution to the SNR occurs as the CO enters the strong field (towards the end of inspiral); due to our initial conditions on $p_{0}$ in this analysis being fixed by the time-to-plunge, the eccentricity at this stage of the inspiral decreases as a function of mass ratio (as the inspiral will circularise more when starting from larger orbital separations). This accentuates the impact of eccentricity on detectability as mass ratio increases, as depicted in the left panel of Fig.˜17.
@@
We also find that the mass location of the detectability peak shows different trends in the two panels, moving towards lower (higher) source frame primary masses with larger mass-ratios (MBH spins). The relationship between mass ratio and SNR (left panel), for a given value of $m_{1}$, is largely driven by the proportionality of the waveform amplitudes to $\mu\sim m_{2}$ (Eq.˜1). As the rate of inspiral also increases with mass ratio, our fixed time-to-plunge leads to higher values of $p_{0}$; as this decreases the frequencies of strong harmonics, this in turn slightly accentuates the detectability of lower-mass systems, which manifests as a gradual shift in the location of the maximum horizon redshift in the left panel of Fig.˜17. The impact of MBH spin (right panel) is also highly pronounced, particularly at larger values of $m_{1}$. Raising MBH spin increases the number of strong-field orbital cycles completed by the CO before it plunges, where both the amplitude and frequency of the GWs emitted are largest. For lower-$m_{1}$ systems, this high-frequency radiation sits towards the upper end of LISA’s sensitivity and therefore does not contribute strongly to the SNR (and therefore the detectability) of these systems. As $m_{1}$ increases, this strong-field emission decreases towards the frequencies where LISA is most sensitive, significantly enhancing the overall SNR. This trend continues until the strong-field emission shifts below the minimum in the LISA sensitivity curve ($\sim 3\,\mathrm{mHz}$; see Fig.˜1) at which point the SNR begins to decrease.
@@ V.5 Distinguishing quasi-circular and mildly eccentric systems
Figure 18: Mismatches (blue dots) between quasi-circular ($h^{\mathrm{circ}}$) and eccentric $h^{\mathrm{ecc}}$ waveforms as $e_{0}\rightarrow 0$. Other source parameters $\{m_{1},m_{2},a\}=\{10^{6}\,M_{\odot},25\,M_{\odot},0.998\}$, with SNR $\rho=100$. The black dashed line indicates the Lindblom distinguishability criterion (Eq.˜15) for this SNR, with $D=12$ (12 sampled parameters). For $e_{0}\gtrsim 10^{-5}$, mismatches grow as $(e_{0})^{4}$ (red dashed line) in line with PN scaling arguments. At lower initial eccentricities, other systematics in waveform generation obfuscate any physical relationship between mismatch and initial eccentricity.
@@
We compute mismatches $\mathcal{M}(h^{\rm circ},h^{\rm ecc})$ for $e_{0}\in[10^{-8},10^{-3}]$. Our results are given in Fig.˜18. For very low eccentricities $e_{0}\in[10^{-8},10^{-5})$, the mismatch between circular and eccentric waveforms is dominated by ODE numerical errors. However, for $e_{0}\gtrsim 10^{-5}$, the mismatch grows by an amount proportional to $(e_{0}^{2})^{2}$ as given by the red dashed line. The black dashed line indicates the Lindblom distinguishability criterion (Eq.˜15), where mismatch values $(<)>D/2\rho^{2}$ indicate that the true and approximate waveforms are (in)distinguishable. As in Section˜IV.3, we set the number of sampled model parameters $D=12$. From Fig.˜18, we see that the two models are distinguishable for $e_{0}\gtrsim 2\cdot 10^{-4}$, highlighting that for $\rho<100$ we cannot resolve eccentricities smaller than $e_{0}<2\cdot 10^{-4}$ for this set of source parameters. We have checked that our conclusions remain unchanged when decreasing the ODE integrator error $\sigma_{\mathrm{tol}}$.
@@ V.6 Inference prospects for asymmetric-mass binaries with LISA
11 EMRI (Prograde) $10^{6}$ $10^{1}$ $0.998$ $7.728$ 0.730(0.045)0.730\,(0.045) $2.204$ $0.394$ $50$
22 Strong-field EMRI $10^{7}$ $10^{1}$ $0.998$ $2.120$ 0.425(0.261)0.425\,(0.261) $3.590$ $0.593$ $30$
33 Heavy IMRI $10^{7}$ $10^{5}$ $0.950$ $23.425$ 0.850(0.023)0.850\,(0.023) $7.250$ $1.058$ $500$
44 Light IMRI $10^{5}$ $10^{3}$ $0.950$ $74.383$ 0.850(0.004)0.850\,(0.004) $3.500$ $0.581$ $200$
55 EMRI (Retrograde) $10^{5}$ $10^{1}$ 0.500-0.500 $26.190$ 0.800(0.195)0.800\,(0.195) $1.081$ $0.212$ $30$
@@
Table 1: Parameters for investigated EMRI and IMRI sources. Waveforms for these systems are shown in Fig.˜19. (From left to right columns): Detector-frame masses $m_{1}$ and $m_{2}$, dimensionless MBH spin $a$, initial semi-latus rectum $p_{0}$, initial (final) eccentricity $e_{0}$ ($e_{\mathrm{f}}$), luminosity distance $d_{\mathrm{L}}$, redshift $z_{\mathrm{r}}$ and the optimal SNR $\rho$. Each source is observed for two years and plunges just after the end of the observation window. For sources with $m_{1}=10^{5}M_{\odot}$, we use a finer sample cadence $\mathrm{d}t=2\,\mathrm{s}$ to resolve high-frequency GWs near plunge; all other sources use our fiducial sampling interval of $\mathrm{d}t=5\,\mathrm{s}$. In all cases, we fix the angular parameters $(\theta_{S},\phi_{S},\theta_{K},\phi_{K})=(0.8,2.2,1.6,1.2)$ and initial orbital phases $(\Phi_{\phi_{0}},\Phi_{r_{0}})=(2.0,3.0)$.
@@
In this subsection, we investigate parameter estimation prospects for a range of IMRI and EMRI sources. As few is capable of generating waveforms in $\sim 100\,\mathrm{ms}$ (Section˜V.1), we are able to perform Bayesian inference studies on a timescale of hours, directly exploring the structure and scale of the posterior distributions of these sources. The motivation for performing these analyses is twofold. First, we seek to verify that our model reproduces features expected of posterior distributions for these systems: in particular, that they are approximately Gaussian, and that parameters intrinsic to the inspiral dynamics (such as spin and eccentricity) are recovered with high precision. Second, it was identified in Ref. Katz:2021yft that the mode content of EMRI waveform models can be reduced significantly without significantly biasing inference results; our analyses will re-examine this conclusion with the inclusion of MBH spin and for IMRI sources. To our knowledge, this is the first full Bayesian investigation of rapidly-spinning and highly eccentric IMRI and EMRI systems to appear in the literature to date (however, see Ref. Huez:2025npe for a parameter estimation study of mildly eccentric binaries with $q\leq 10$).
@@
Figure 19: Time-domain snapshots of the science-case waveforms described in Table˜1 during the early (left panel) and late (right panel) periods of a two-year inspiral. This set of representative sources demonstrates the wide range of waveform morphologies captured by our model. In all the panels, only the plus polarization $h_{+}$ is shown; for the two IMRI systems, initial orbital phases have also been adjusted for visual clarity.
@@
We perform parameter estimation on simulations of each of the sources listed in Table˜1. The parameters of these sources were chosen to represent the wide range of waveform morphologies accessible with our waveform model Fig.˜19, with luminosity distances set to astrophysically-motivated values. A notable exception to our analysis conventions is that we set a finer sampling cadence of $\mathrm{d}t=2\,\mathrm{s}$ for Sources 4 and 5 in order to better resolve their high-frequency mode content; for other sources, we use our fiducial sampling cadence of $\mathrm{d}t=5\,\mathrm{s}$. While these sampling cadences are not quite high enough to resolve the highest-frequency harmonics near plunge for some of these systems (particularly Sources 1 and 4), we found that these cadences were sufficient to obtain accurate parameter constraints. For each simulation, we generate a waveform with mode-selection parameter $\kappa=0$ (i.e., all modes in our amplitude model are included). During inference, we consider $\kappa=10^{-5}$ and $\kappa=\{10^{-3},10^{-2}\}$ (the latter value depending on the SNR of the injection), demonstrating how parameter inference results are impacted when weaker harmonics (which are more important for high eccentricities and spins; see Fig.˜4) are neglected. We do not show the (marginal) posteriors obtained from our inference runs in the main text for brevity; they can be found in Appendix˜F, and are referenced in the right-most column of Tab. 2, where we summarize the relative measurement precisions for parameters of astrophysical interest. These include the source frame masses $m_{i}^{(\rm s)}$, spin $a$, final eccentricity $e_{f}$ and luminosity distance $d_{\mathrm{L}}$. We also report the sky-localization area of each source: given the covariance matrix of the sky-angle posterior samples $\Sigma=\rm cov(\theta_{S},\,\phi_{S})$, and $\bar{\theta}_{S}=\rm{median}(\theta_{S})$, we compute the $99\%$ quantile of the sky-localization area in square degrees, given by